An asset used in a four-year project falls in the five-year MACRS class (MACRS Table) for tax purposes. The asset has an acquisition cost of $7,700,000 and will be sold for $1,820,000 at the end of the project. If the tax rate is 24 percent, what is the aftertax salvage value of the asset? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
Answer #1

Asset Acquisition Cost = $ 7700000, Sale Price = S = $ 1820000, Tax Rate = t = 24 %, Project LIfe = 4 years, Depreciation Strategy: 5-year MACRS Class,

5-Year MACRS Depreciation %: Year 1 = 20%, Year 2 = 32 %, Year 3 = 19.2 %, Year 4 = 11.52 %, Year 5 = 11.52 %, Year 6 = 5.76 %

Year 1: Beginning Value = $ 7700000

Depreciation Expense = 0.2 x 7700000 = $ 1540000

Ending Value = 7700000 - 1540000 = $ 6160000

Year 2: Beginning Value = $ 6160000

Depreciation Expense = 0.32 x 7700000 = $ 2464000

Ending Value = 6160000 - 2464000 = $ 3696000

Year 3: Beginning Value = $ 3696000

Depreciation Expense = 0.192 x 7700000 = $ 1478400

Ending Value = 3696000 - 1478400 = $ 2217600

Year 4: Beginning Value = $ 2217600

Depreciation Expense = 0.1152 x 7700000 = $ 887040

Ending Value = 2217600 - 887040 = $ 1330560

Book Value at Project culmination = Ending Value at the end of Year 4 = BV4 = $ 1330560

After-Tax Salvage Value = Sale Price - Tax Rate (Sale Price - BV4) = 1820000 - 0.24 x (1820000 - 1330560) = $ 1702534